OpenStudy (samigupta8):

The value of lim e^xcosx-1-x/sin(x^2) where x----->0 is equal to..

OpenStudy (samigupta8):

My ans is 0..

OpenStudy (samigupta8):

@hartnn

OpenStudy (samigupta8):

And 1/2 also...how is it possible??

OpenStudy (samigupta8):

I did it by two methods resulting in two diff ans ..

OpenStudy (samigupta8):

First of all i multiplied e^xcosx-1 by xcosx and divided it by same.. Then i got xcosx-x/x^2 Solving this limit i got 0..

hartnn (hartnn):

you have used lim x->0 (e^x-1)/x =1 in the numerator...but is that legal step ? lim (ab+c)/d does not equal lim [(lim a)b+c]/d

OpenStudy (samigupta8):

Can't we write like this Lim (e^xcosx -1 -x ) = lim (e^xcosx-1)- limx

hartnn (hartnn):